Suppose that there are many stocks in the security market and that the characteristics of stocks A and B are given as follows:

Stock Expected Return Standard Deviation
A 14 % 7 %
B 18 9
Correlation = –1

Suppose that it is possible to borrow at the risk-free rate, rf. What must be the value of the risk-free rate? (Hint: Think about constructing a risk-free portfolio from stocks A and B.) (Do not round intermediate calculations. Round your answer to 3 decimal places.)

Risk Free Rate? %
Answer #1

Given,

Stock Expected Return Standard Deviation
A 14% 7%
B 18% 9%
Correlation = –1

Since Correlation between two stock is -1, creating a risk free portfolio is possible using weights

WA = SDB/(SDA + SDB)

So, WA = 9/(7+9) = 0.5625

and WB = 1 - WA = 1-0.5625 = 0.4375

So, risk free rate = weighted average return of stocks = 0.5625*14 + 0.4375*18 = 15.75%

value of the risk-free rate = 15.75%